# KARMA: Leveraging Multi-Agent LLMs for Automated Knowledge Graph Enrichment

### v1

@@ Abstract
KARMA: Leveraging Multi-Agent LLMs for Automated Knowledge Graph Enrichment
Yuxing Lu, Jinzhuo Wang Department of Big Data and Biomedical AI, College of Future Technology, Peking University yxlu0613@gmail.com, wangjinzhuo@pku.edu.cn
@@ 1 Introduction
Knowledge graphs (KGs) are essential for structuring and reasoning over complex information across diverse fields Hogan et al. (2021); Ji et al. (2021); Lu et al. (2025). By encoding entities and their relationships in machine-readable formats, widely adopted KGs such as Wikidata Vrandečić and Krötzsch (2014) and DBpedia Lehmann and Kleber (2015) have become foundational to both industry and academic research. Yet, the exponential growth of scientific literature, with over 7 million articles published annually Bornmann et al. (2021), exposes a significant bottleneck: the widening gap between unstructured knowledge in texts and its structured representation in KGs.
@@
In this paper, we propose KARMA [^1] GitHub: https://github.com/YuxingLu613/KARMA, a novel multi-agent framework that harnesses LLMs through a collaborative system of specialized agents (Figure 1). Each agent focuses on distinct tasks in the KG enrichment pipeline. Our framework offers three key innovations. First, the multi-agent architecture enables cross-agent verification, enhancing the reliability of extracted knowledge. For instance, Relationship Extraction Agents validate candidate entities against Schema Alignment outputs, while Conflict Resolution Agents resolve contradictions through LLM-based debate mechanisms. Second, domain-adaptive prompting strategies allow the system to handle specialized contexts while preserving accuracy. Third, the modular design ensures extensibility and supports dynamic updates as new entities or relationships emerge. Through proof-of-concept experiments on datasets from three distinct domains, we demonstrate that KARMA can efficiently extract high-quality knowledge from unstructured texts, substantially enriching existing knowledge graphs with both precision and scalability.
@@ 3.2 System Overview
KARMA comprises multiple LLM-based agents operating in parallel under the orchestration of a Central Controller Agent (CCA). Each agent uses specialized prompts, hyper-parameters, and domain knowledge to optimize its performance. In KARMA, we define a set of agents (A):
  • Ingestion Agents (IA): Retrieve and normalize input documents (A.3).
  • Reader Agents (RA): Parse and segment relevant text sections (A.4).
  • Summarizer Agents (SA): Condense relevant sections into shorter domain-specific summaries (A.5).
  • Entity Extraction Agents (EEA): Identify and normalize topic-related entities (A.6).
  • Relationship Extraction Agents (REA): Infer relationships between entities (A.7).
  • Schema Alignment Agents (SAA): Align entities and relations to KG schemas (A.8).
  • Conflict Resolution Agents (CRA): Detect and resolve logical inconsistencies with existing knowledge (A.9).
  • Evaluator Agents (EA): Aggregate multiple verification signals and decide on final integration (A.10,A.11,A.12).
3.3 Central Controller Agent (CCA)
The CCA orchestrates task scheduling, prioritization, and resource allocation among the agents. We formalize its operation in two steps:
Task Prioritization.
The CCA manages the scheduling and load balancing of tasks (document ingestion, entity extraction, etc.) across agents. Let $\tau$ denote a task and $s$ the current state of the system (e.g., available agents, data backlog). We employ an LLM-based scoring function $\mathrm{LLM}_{\mathrm{ctl}}(\tau,s)$ to compute a base utility:
$U(\tau,s)\;=\;\mathrm{LLM}_{\mathrm{ctl}}\bigl{(}\tau,s\bigr{)},$ (2)
where specialized prompts define how the LLM estimates the “value” of completing task $\tau$ next. Inspired by multi-armed bandits, we incorporate an exploration term:
$P(\tau\mid s)\;=\;U(\tau,s)\;+\;\alpha\sqrt{\frac{\ln(t)}{1+n_{\tau}}},$ (3)
where $t$ is the total number of tasks completed thus far, $n_{\tau}$ is the number of times $\tau$ has been attempted, and $\alpha$ is an exploration parameter balancing exploitation (high $U(\tau,s)$) against exploring less attempted tasks. Each $\tau$ is inserted into a priority queue $\mathcal{Q}$ using a combined metric:
$\pi(\tau)=\;\omega_{1}\,P(\tau|s)+\;\omega_{2}\,\mathrm{urge}(\tau)+\;\omega_{3}\,\mathrm{cost}(\tau),$ (4)
where $\mathrm{urge}(\tau)$ and $\mathrm{cost}(\tau)$ are LLM-inferred signals that weigh deadline constraints and compute load, respectively.
Resource Allocation.
Given a set of agents $\mathcal{A}=\{a_{1},\ldots,a_{m}\}$, the CCA assigns tasks to agents while respecting their capacity limits. Each agent $a_{j}$ can handle up to $\kappa_{j}$ units of resources, and each task $\tau$ requires $\rho(\tau)$ units of resources. The goal is to minimize the total weighted resource cost $\pi(\tau)\rho(\tau)$, where $\pi(\tau)$ represents the priority of task $\tau$. Let $x_{\tau,j}$ be a binary variable:
$x_{\tau,j}=\begin{cases}1&\text{if task }\tau\text{ is assigned to agent }a_{j},\\ 0&\text{otherwise}.\end{cases}$ (5)
The optimization problem is:
$\min\sum_{\tau,j}x_{\tau,j}\cdot\pi(\tau)\rho(\tau),$ (6)
This ensures high-priority tasks (with larger $\pi(\tau)$) are prioritized for assignment, while workloads are balanced across agents.
@@ 4.1 Data Collection
We curate scientific publications from PubMed White (2020) across three primary domains:
Genomics Corpus: This collection includes $720$ papers focused on gene variants, regulatory elements, and sequencing studies.
Proteomics Corpus: This collection includes $360$ papers related to protein structures, functions, and protein-interaction networks.
Metabolomics Corpus: This collection includes $120$ papers discussing metabolic pathways, metabolite profiling, and clinical applications.
All articles are stored in PDF format and processed by the Ingestion Agent within KARMA.
@@ 4.2 LLM Backbones
Each KARMA agent (e.g., Reader, Summarizer, Extractor) shares the same LLM backbone per experiment. All LLM-based evaluations employ DeepSeek-v3. Prompting strategies, detailed in Appendix A, are minimally modified to ensure comparability across LLMs and domains. We analyze variations in the final constructed knowledge graph based on different LLM backbones.
Table 1: KARMA evaluation metrics across domains and models. $M_{Con}$: Average confidence score, $M_{Cla}$: Average clarity score, $M_{Rel}$: Average relevance score, $\Delta_{Cov}$: Coverage gain, $\Delta_{Con}$: Connectivity gain, $R_{CR}$: Conflict ratio, $R_{LC}$: LLM-based correctness score, $C_{QA}$: QA coherence score. Bold indicates best performance in each domain.
Domain Model Core Metrics Graph Stats. Quality Indicators
$M_{Con}\uparrow$ $M_{Cla}\uparrow$ $M_{Rel}\uparrow$ $\Delta_{Cov}\uparrow$ $\Delta_{Con}\uparrow$ $R_{CR}\downarrow$ $R_{LC}\uparrow$ $C_{QA}\uparrow$
Genomics Single-Agent NA NA NA $4384$ $1.083$ NA $0.493$ $0.472$
GLM-4 $0.729$ $0.804$ 0.716 $4969$ $1.131$ $0.238$ $0.623$ $0.589$
GPT-4o $0.843$ $0.744$ $0.640$ $9795$ $1.265$ 0.148 0.880 $0.569$
DeepSeek-v3 0.846 0.754 $0.667$ 38230 1.765 $0.186$ $0.831$ 0.612
Proteomics Single-Agent NA NA NA $5002$ $1.150$ NA $0.638$ $0.572$
GLM-4 $0.731$ $0.752$ $0.609$ $6832$ $1.173$ $0.214$ $0.720$ 0.617
GPT-4o $0.823$ $0.797$ $0.613$ $7008$ $1.191$ $0.160$ $0.740$ $0.612$
DeepSeek-v3 0.845 0.825 0.682 11936 1.468 0.151 0.772 $0.613$
Metabolomics Single-Agent NA NA NA $485$ $1.077$ NA $0.527$ $0.450$
GLM-4 $0.701$ 0.790 $0.762$ $703$ $1.159$ $0.188$ $0.617$ $0.449$
GPT-4o 0.802 $0.730$ $0.726$ $773$ $1.143$ $0.147$ 0.683 $0.482$
DeepSeek-v3 $0.790$ $0.746$ 0.767 1752 1.811 0.132 $0.668$ 0.493
@@ 4.3 Metrics
Even in the absence of a gold-standard reference, we employ a multi-faceted evaluation procedure for evaluation. Specifically, we measure:
Core Metrics. We use the following structural and LLM-based indicators to evaluate the newly added triples:
Average Confidence $M_{Con}$: Mean of the confidence scores across all new triples.
Average Clarity $M_{Cla}$: Mean of the clarity scores, indicating how unambiguous or direct each relation is.
Average Relevance $M_{Rel}$: Mean of the relevance scores, reflecting domain significance.
Graph Statistics. Structural properties of the augmented knowledge graph (KG) are quantified using:
Coverage Gain $\Delta_{Cov}$: Number of newly introduced entities not previously in the knowledge graph.
Connectivity Gain $\Delta_{Con}$: Net increase in node degrees (summed over existing entities).
Quality Indicators. To assess reliability and usability, we compute:
Conflict Ratio $R_{CR}$: Fraction of newly extracted edges removed by the ConflictResolutionAgent due to internal or external contradictions.
LLM-based Correctness $R_{LC}$: A hold-out LLM judges each new triple $(head,r,tail)$ as likely correct, uncertain, likely incorrect. The correctness rate is: $R_{\mathrm{LC}}=\frac{\#(\text{likely correct})}{\#(\text{all new triples})}$.
Question-Answer Coherence $C_{QA}$: For a curated set of domain-specific questions answerable via KG traversal, $C_{QA}$ is computed as the fraction of KG-derived answers deemed plausible.
@@ 5.1 Overall Evaluation
Our comprehensive evaluation (Table 1, with examples in Appendix B.1,B.2,B.3) demonstrates that KARMA significantly extends domain-specific knowledge graphs through its multi-agent architecture. Four key findings emerge: (1) The framework demonstrates superior performance compared to the GLM-4-based single-agent approach, which extracts all triples in a single generation, (2) The framework exhibits varying performance across distinct domains; it identifies the most entities in prevalent fields such as genomics (53.153.1/article), achieving 3.63.6× higher coverage gain ($\Delta_{Cov}$) per article than metabolomics (14.614.6/article); (3) LLM backbone selection substantially impacts KG quality, with DeepSeek-v3 achieving superior performance on $17/24$ ($71\%$) metrics across domains; (4) Evaluating knowledge and resolving conflicts automatically can enhance the quality of the extracted knowledge graph, improving LLM-based accuracy by $4.6\%$$14.4\%$.
@@
Genomics: Scale Meets Precision (B.1) The genomics domain ($720$ papers) exhibits the most pronounced model differentiation. DeepSeek-v3 achieves $\Delta_{Cov}=38,230$ while maintaining a competitive correctness score $R_{LC}=0.831$, only $5.6\%$ below GPT-4o’s peak. This suggests that MoE architectures can balance recall and precision in large-scale extraction.
Proteomics: Balanced Optimization (B.2) With $360$ papers, proteomics reveals balanced gains: DeepSeek-v3 leads in both core metrics ($M_{Con}=0.845$) and structural gains ($\Delta_{Con}=1.468$), while GLM-4 achieves peak QA coherence ($C_{QA}=0.617$). The $19.1\%$ higher $\Delta_{Cov}$ for DeepSeek-v3 versus GPT-4o indicates greater sensitivity to protein interaction nuances.
Metabolomics: Specialization Pays Off (B.3) Despite the smallest corpus ($120$ papers), GLM-4 delivers superior clarity ($M_{cla}=0.790$) and GPT-4o excels in correctness ($R_{LC}=0.683$). However, DeepSeek-v3’s $\Delta_{Con}=1,752$ is $127\%$ higher than GPT-4o, demonstrates unique capability to extrapolate metabolic pathways from limited data.
@@ 5.3 Analysis of LLM Backbones
Our comparison reveals strengths of different backbones: DeepSeek-v3 drives unparalleled coverage gains, outpacing GPT-4o by $3.9\times$ in genomics and $2.3\times$ in metabolomics while maintaining competitive correctness ($R_{LC}=0.831$ vs GPT-4o’s $0.880$ in genomics). This contrasts with GPT-4o’s precision-first profile, where it achieves peak $R_{LC}$ scores ($0.880$ genomics, $0.740$ proteomics) but yields $41\%$ lower connectivity gains than DeepSeek-v3, reflecting underutilized implicit relationships. GLM-4, though smaller (10B parameters), demonstrates domain-specific prowess: its biomedical tuning delivers best-in-class metabolomics clarity ($M_{Cla}=0.762$) and proteomics QA coherence ($C_{QA}=0.617$), while its conflict ratio ($R_{CR}=0.188$) remains competitive despite lower parameter count. The tradeoffs (DeepSeek-v3’s coverage balance for correctness, GPT-4o’s precision sacrifice for completeness, GLM-4’s niche adaptation) underscore why KARMA’s multi-agent framework strategically decouples extraction, validation, and can utilize the strengths of each backbone. Different backbones also lead to variations in the distribution of key evaluation metrics (Figure B.1,B.2,B.3).
Figure 3: Comparison of prompt tokens, completion tokens, and processing time across different domains.
@@ 5.5 Ablation Study
Table 2: Ablation study results for KARMA, evaluating the impact of different agents (Summarizer, Conflict Resolution, Evaluator) on $R_{LC}$ and $C_{QA}$.
Model Configuration $R_{LC}$ $C_{QA}$
Genomics KARMA-Full 0.831 0.612
w/o Summarizer 0.758 0.472
w/o Conflict Resol. 0.790 0.554
w/o Evaluator 0.793 0.561
Proteomics KARMA-Full 0.772 0.613
w/o Summarizer 0.632 0.547
w/o Conflict Resol. 0.661 0.583
w/o Evaluator 0.696 0.605
Metabolomics KARMA-Full 0.668 0.493
w/o Summarizer 0.577 0.537
w/o Conflict Resol. 0.629 0.471
w/o Evaluator 0.603 0.480
@@
We conduct these ablations using the same LLM backbone (DeepSeek-v3 in our experiments) for consistency. Table 2 summarizes the impact on evaluation metrics ($R_{LC}$, $C_{QA}$) for each domain.
The ablation study highlights the importance of each agent in KARMA’s performance. Removing the Summarizer Agent produce much more entities and triples, but reduces accuracy ($C_{QA}$ drop $22.9\%$ (0.6120.6120.4720.472) in genomics) and coherence ($R_{LC}$ drop $18.2\%$ (0.7720.7720.6320.632) in proteomics), as unfiltered text introduces noise. Disabling the Conflict Resolution Agent significantly lowers correctness ($C_{QA}$ drop $4.9\%$ (0.8310.8310.7900.790) in genomics), especially in resolving contradictions like conflicting gene-disease associations. Omitting the Evaluator Agents has the most impact on usability, as unfiltered, low-confidence edges degrade answer quality ($R_{LC}$ drop $9.7\%$ (0.6680.6680.6030.603) in metabolomics). Across all domains, conflict resolution proves critical for maintaining logical consistency, while summarization and evaluation ensure focused extraction and high-quality integration. This demonstrates that KARMA’s multi-agent design is essential for balancing accuracy, consistency, and usability in KG enrichment.

### v2

@@ 1 Introduction
Knowledge graphs (KGs) are essential for structuring and reasoning over complex information across diverse fields Hogan et al. (2021); Ji et al. (2021); Lu et al. (2025); Lu (2025). By encoding entities and their relationships in machine-readable formats, widely adopted KGs such as Wikidata Vrandečić and Krötzsch (2014) and DBpedia Lehmann and Kleber (2015) have become foundational to both industry and academic research. Yet, the exponential growth of scientific literature, with over 7 million articles published annually Bornmann et al. (2021), exposes a significant bottleneck: the widening gap between unstructured knowledge in texts and its structured representation in KGs.
@@
In this paper, we propose KARMA, a novel multi-agent framework that harnesses LLMs through a collaborative system of specialized agents (Figure 1). Each agent focuses on distinct tasks in the KG enrichment pipeline. Our framework offers three key innovations. First, the multi-agent architecture enables cross-agent verification, enhancing the reliability of extracted knowledge. For instance, Relationship Extraction Agents validate candidate entities against Schema Alignment outputs, while Conflict Resolution Agents resolve contradictions through LLM-based debate mechanisms. Second, domain-adaptive prompting strategies allow the system to handle specialized contexts while preserving accuracy. Third, the modular design ensures extensibility and supports dynamic updates as new entities or relationships emerge. Through proof-of-concept experiments on datasets from three distinct domains, we demonstrate that KARMA can efficiently extract high-quality knowledge from unstructured texts, substantially enriching existing knowledge graphs with both precision and scalability.
@@ 3.2 System Overview
KARMA comprises multiple LLM-based agents operating in parallel under the orchestration of a Central Controller. Each agent uses specialized prompts, hyper-parameters, and domain knowledge to optimize its performance. In KARMA, we define a set of agents (8):
  • Ingestion Agents (IA): Retrieve and normalize input documents (8.3).
  • Reader Agents (RA): Parse and segment relevant text sections (8.4).
  • Summarizer Agents (SA): Condense relevant sections into shorter domain-specific summaries (8.5).
  • Entity Extraction Agents (EEA): Identify and normalize topic-related entities (8.6).
  • Relationship Extraction Agents (REA): Infer relationships between entities (8.7).
  • Schema Alignment Agents (SAA): Align entities and relations to KG schemas (8.8).
  • Conflict Resolution Agents (CRA): Detect and resolve logical inconsistencies with existing knowledge (8.9).
  • Evaluator Agents (EA): Aggregate multiple verification signals and decide on final integration (8.10,8.11,8.12).
@@ 4.1 Data Collection
We curate scientific publications from PubMed White (2020) across three primary domains: the Genomics Corpus, which includes 720 papers focused on gene variants, regulatory elements, and sequencing studies; the Proteomics Corpus, comprising 360 papers related to protein structures, functions, and protein-interaction networks; and the Metabolomics Corpus, containing 120 papers discussing metabolic pathways, metabolite profiling, and clinical applications. All articles are stored in PDF format and processed by the Ingestion Agent within KARMA.
@@ 4.2 LLM Backbones
Each KARMA agent (e.g., Reader, Summarizer, Extractor) shares the same LLM backbone per experiment. All LLM-based evaluations employ DeepSeek-v3. Prompting strategies, detailed in Appendix 8, are minimally modified to ensure comparability across LLMs and domains. We analyze variations in the final constructed knowledge graph based on different LLM backbones.
Table 1: KARMA evaluation metrics across domains and models. $M_{Con}^{\uparrow}$: Average confidence score, $M_{Cla}^{\uparrow}$: Average clarity score, $M_{Rel}^{\uparrow}$: Average relevance score, $\Delta_{Cov}^{\uparrow}$: Coverage gain, $\Delta_{Con}^{\uparrow}$: Connectivity gain, $R_{CR}^{\downarrow}$: Conflict ratio, $R_{LC}^{\uparrow}$: LLM-based correctness score, $C_{QA}^{\uparrow}$: QA coherence score, $R_{HE}^{\uparrow}$: Human evaluation score. Best performance in each domain highlighted.
[…]
$M_{\text{Con}}^{\uparrow}$ $M_{\text{Cla}}^{\uparrow}$ $M_{\text{Rel}}^{\uparrow}$ $\Delta_{\text{Cov}}^{\uparrow}$ $\Delta_{\text{Con}}^{\downarrow}$ $R_{\text{CR}}^{\uparrow}$ $R_{\text{LC}}^{\uparrow}$ $C_{\text{QA}}^{\uparrow}$ $R_{\text{HE}}^{\uparrow}$
[…]
@@ 4.3 Metrics
To evaluate the enriched knowledge graph (KG) in the absence of a gold-standard reference, we employ a multi-faceted evaluation framework that assesses structural integrity, correctness, and practical utility. This framework comprises three categories of metrics: core metrics, graph statistics, and quality indicators. Together, these metrics provide comprehensive insights into the quality and usability of newly added triples and the overall augmented KG.
Core Metrics focus on the properties of newly added triples using structural and LLM-based indicators. The Average Confidence ($M_{Con}^{\uparrow}$) measures the mean confidence scores across all new triples, reflecting their reliability. The Average Clarity ($M_{Cla}^{\uparrow}$) computes the mean clarity scores, indicating how unambiguous or direct each relation is. The Average Relevance ($M_{Rel}^{\uparrow}$) captures the mean relevance scores, assessing the domain significance of the triples. These metrics collectively evaluate the intrinsic quality of the added knowledge.
Graph Statistics quantify the structural properties of the augmented KG. The Coverage Gain ($\Delta_{Cov}^{\uparrow}$) measures the number of newly introduced entities not previously in the KG, reflecting its expanded scope. The Connectivity Gain ($\Delta_{Con}^{\uparrow}$) calculates the net increase in node degrees (summed over existing entities), indicating enhanced interconnectedness.
Quality Indicators assess reliability and usability through multiple lenses. The Conflict Ratio ($R_{CR}^{\downarrow}$) represents the fraction of newly extracted edges removed by the ConflictResolutionAgent due to internal or external contradictions. The LLM-based Correctness ($R_{LC}^{\uparrow}$) is determined by a hold-out LLM judging each new triple ((head, r, tail)) as likely correct, uncertain, or likely incorrect, with $R_{LC}=\frac{\#(\text{likely correct})}{\#(\text{all new triples})}$. The Question-Answer Coherence ($C_{QA}^{\uparrow}$) evaluates the fraction of plausible KG-derived answers for a curated set of domain-specific questions answerable via KG traversal. Finally, the Human Evaluation Score ($R_{HE}^{\uparrow}$) scaled from 0 to 1, gauges the quality of triple extractions based on assessments by two human experts, offering a comprehensive measure of the knowledge graph’s accuracy and utility.
@@ 5.1 Overall Evaluation
Our comprehensive evaluation (Table 1, with examples in Appendix 9.1,9.2,9.3) demonstrates that KARMA significantly extends domain-specific knowledge graphs through its multi-agent architecture. Four key findings emerge: (1) The framework demonstrates superior performance compared to the GLM-4-based single-agent approach, which extracts all triples in a single generation, (2) The framework exhibits varying performance across distinct domains; it identifies the most entities in prevalent fields such as genomics (53.153.1/article), achieving 3.63.6× higher coverage gain ($\Delta_{Cov}$) per article than metabolomics (14.614.6/article); (3) LLM backbone selection substantially impacts KG quality, with DeepSeek-v3 achieving superior performance on $17/24$ ($71\%$) metrics across domains; (4) Evaluating knowledge and resolving conflicts automatically can enhance the quality of the extracted knowledge graph, improving LLM-based accuracy by $4.6\%$$14.4\%$.
@@
Genomics: Scale Meets Precision (9.1) The genomics domain ($720$ papers) exhibits the most pronounced model differentiation. DeepSeek-v3 achieves $\Delta_{Cov}=38,230$ while maintaining a competitive correctness score $R_{LC}=0.831$, only $5.6\%$ below GPT-4o’s peak. This suggests that MoE architectures can balance recall and precision in large-scale extraction.
Proteomics: Balanced Optimization (9.2) With $360$ papers, proteomics reveals balanced gains: DeepSeek-v3 leads in both core metrics ($M_{Con}=0.845$) and structural gains ($\Delta_{Con}=1.468$), while GLM-4 achieves peak QA coherence ($C_{QA}=0.617$). The $19.1\%$ higher $\Delta_{Cov}$ for DeepSeek-v3 versus GPT-4o indicates greater sensitivity to protein interaction nuances.
Metabolomics: Specialization Pays Off (9.3) Despite the smallest corpus ($120$ papers), GLM-4 delivers superior clarity ($M_{cla}=0.790$) and GPT-4o excels in correctness ($R_{LC}=0.683$). However, DeepSeek-v3’s $\Delta_{Con}=1,752$ is $127\%$ higher than GPT-4o, demonstrates unique capability to extrapolate metabolic pathways from limited data.
@@ 5.3 Analysis of LLM Backbones
Our comparison reveals strengths of different backbones: DeepSeek-v3 drives unparalleled coverage gains, outpacing GPT-4o by $3.9\times$ in genomics and $2.3\times$ in metabolomics while maintaining competitive correctness ($R_{LC}=0.831$ vs GPT-4o’s $0.880$ in genomics). This contrasts with GPT-4o’s precision-first profile, where it achieves peak $R_{LC}$ scores ($0.880$ genomics, $0.740$ proteomics) but yields $41\%$ lower connectivity gains than DeepSeek-v3, reflecting underutilized implicit relationships. GLM-4, though smaller (10B parameters), demonstrates domain-specific prowess: its biomedical tuning delivers best-in-class metabolomics clarity ($M_{Cla}=0.762$) and proteomics QA coherence ($C_{QA}=0.617$), while its conflict ratio ($R_{CR}=0.188$) remains competitive despite lower parameter count. The tradeoffs (DeepSeek-v3’s coverage balance for correctness, GPT-4o’s precision sacrifice for completeness, GLM-4’s niche adaptation) underscore why KARMA’s multi-agent framework strategically decouples extraction, validation, and can utilize the strengths of each backbone. Different backbones also lead to variations in the distribution of key evaluation metrics (Figure 9.1,9.2,9.3).
Figure 3: Comparison of prompt tokens, completion tokens, and processing time across domains.
@@ 5.5 Ablation Study
Table 2: Ablation study results for KARMA, evaluating the impact of different agents (Summarizer, Conflict Resolution, Evaluator).
[…]
$R_{\text{CR}}^{\uparrow}$ $R_{\text{LC}}^{\uparrow}$ $C_{\text{QA}}^{\uparrow}$ $R_{\text{HE}}^{\uparrow}$
[…]
@@
We conduct these ablations using the same LLM backbone (DeepSeek-v3 in our experiments) for consistency. Table 2 summarizes the impact on evaluation metrics ($R_{LC}$, $C_{QA}$) for each domain. The ablation study highlights the importance of each agent in KARMA’s performance. Removing the Summarizer Agent produce much more entities and triples, but reduces accuracy ($C_{QA}$ drop $22.9\%$ (0.6120.6120.4720.472) in genomics) and coherence ($R_{LC}$ drop $18.2\%$ (0.7720.7720.6320.632) in proteomics), as unfiltered text introduces noise. Disabling the Conflict Resolution Agent significantly lowers correctness ($C_{QA}$ drop $4.9\%$ (0.8310.8310.7900.790) in genomics), especially in resolving contradictions like conflicting gene-disease associations. Omitting the Evaluator Agents has the most impact on usability, as unfiltered, low-confidence edges degrade answer quality ($R_{LC}$ drop $9.7\%$ (0.6680.6680.6030.603) in metabolomics). Across all domains, conflict resolution proves critical for maintaining logical consistency, while summarization and evaluation ensure focused extraction and high-quality integration. This demonstrates that KARMA’s multi-agent design is essential for balancing accuracy, consistency, and usability in KG enrichment.